Verify if the given value of x is a solution to the equation.
1. 2x - 3 = 9, x=-6
2.-3(x + 5)= 2x - 11, x=-4/5

Answer by Alan3354(69443)

Verif30 " y if the given value of x is a solution to the equation.
1. 2x - 3 = 9, x=-6
2.-3(x + 5)= 2x - 11, x=-4/5
----------------------------
Sub the given value for x and see if the 2 sides are equal.
----
eg,
3x - 15 = 0 : x = 5
3*5 - 15 = 0
15 - 15 = 0
0 = 0
---------------------------
4x - 2 = 30 : x = 6
4*6 - 2 = 30
24 - 2 = 30
22 = 30 >>>>>>>>>>> Not a solution